Using the following information, your maximum profits (positive) would be $____.

  • Spot rate: $0.5/NZ$
  • Expectation: ZN$ will appreciate depreciate against $ by 6%
  • You can borrow either $5,000 or NZ$10,000
  • Given interest rates:

Lending rate

Borrowing rate

US$

6.8% (APR)

7.1%(APR)

NZ$

6.6%(APR)

6.9%(APR)
Answer #1

Case1: Assume $ NZ appreciates 6%

I will borrow $5000 at 7.1%

convert to $NZ and I will get NZ$10000

will lend NZ$10000 at 6.6%

After the time period will get back 10000*1.066=NZ$10660

will convert NZ$10660 to $-------> 10660*0.5*1.06(since NZ$ appreciates 6%)---->$5649.8

The amount to be paid back including interest=$5000*(1+7.1%)=5355

Net profit=5649.8-5355=$294.8

Case2:

Assume $ appreciates 6% wrt NZ$

I will borrow NZ$10000 at 6.9%

convert to $ to NZ$ and I will get $5000

will lend $5000 at 6.8%

After the time period will get back 5000*1.068=$5340

will convert $5340 to NZ$-------> 5340*2*1.06(since $ appreciates 6%)---->NZ$11320.8

The amount to be paid back including interest=NZ$10000*(1+6.9%)=NZ$10690

Net profit=11320-10690=NZ$630.8

lets convert this to $=630.8*0.5/1.06=$297.55 This is the max profit

Hit like if you like my answer